A first order reaction, where [ A ] o = 1.3 M, is 70.9 % complete in 324 s. How long does it take for the same reaction to go from 1.3 M to 87.9 % completion?
Answer #1

100 1.3 Ans Given [Ao]=Q = 1.3 M t> 324 S. ; X= 70.7%. (Q-x) = 100-70.9% = 29.1% Rate constant k=? r = 1.3x29.1. ke Hladal (